OpenStudy (lollylau):

James and Tom run around a track of 4200 m. They start together. James runs at a speed at 140m/minute. Tom runs at 100m/s. James rests 1 minute when he goes back to the start. Tom rests 3 minutes when he goes back to the start. When do they meet again? And how do I do questions like this?

OpenStudy (king):

they ran around a track of 4200 m what?radius ,circumference,area?

OpenStudy (lollylau):

circumference

OpenStudy (king):

they rest for a certain time and start again rite?

OpenStudy (lollylau):

yep

OpenStudy (king):

first convert 140m/min to m/s =140*1/60=7/3

OpenStudy (lollylau):

tom's speed is 100/min, but never mind

OpenStudy (king):

what !!tom;s speed is 100m/min or s?

OpenStudy (lollylau):

MINUTE! =.=

OpenStudy (king):

ok so find time taken for both to complete one round of the circle james=4200/140=30 minutes tom=4200/100=42 minutes

OpenStudy (lollylau):

ok

OpenStudy (king):

so if they meet after x rounds of the circle the time taken by each of them will be 30x and 42x

OpenStudy (lollylau):

but then they rest rite?

OpenStudy (king):

oh yeah sry it will be 31x and 45x[30x+x,42x+3x]

OpenStudy (lollylau):

then...

OpenStudy (phi):

The easiest way to do this is find the average speed, S= D/T where T includes the rest time. Then find S1*T mod 1400 = S2*T mod 1400 (this assumes they meet somewhere on the track. Then, to be sure, check that they do not meet at the start while one or the other is resting before they meet on the track.